Dublin University Rifle Club

Dublin University Rifle Club ·Trinity College Dublin ·Dublin 2 ·Ireland

Mark Dennehy 167 Applewood Hts Greystones Co. Wicklow Ireland

Mark Dennehy,

DURC Match Secretary

☎: +353-87-2258207

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2000, current scores  9/12/2000

A Chara,

TTT scores have not been going out on schedule as was originally planned, mainly due to your author getting a lot more work than was originally planned ! Here are the results to date, up to Round 6.

Hopefully I can keep the records a lot more current for the remainder of the competition. Again, remember that the results are also posted on the DURC website : http://www.tcd.ie/clubs/rifle

Mise le meas,

Dublin University Rifle Club ·Trinity College Dublin ·Dublin 2 ·Ireland

Dave Cooney Ardnaglass Corrig Road Dun Laoghaire

Mark Dennehy,

DURC Match Secretary

☎: +353-87-2258207

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2000, current scores  9/12/2000

A Chara,

TTT scores have not been going out on schedule as was originally planned, mainly due to your author getting a lot more work than was originally planned ! Here are the results to date, up to Round 6.

Hopefully I can keep the records a lot more current for the remainder of the competition. Again, remember that the results are also posted on the DURC website : http://www.tcd.ie/clubs/rifle

Mise le meas,

Dublin University Rifle Club ·Trinity College Dublin ·Dublin 2 ·Ireland

John K. Walsh Carjon Lodge Cross Chapel Blessington Co. Wicklow

Mark Dennehy,

DURC Match Secretary

☎: +353-87-2258207

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2000, current scores  9/12/2000

A Chara,

TTT scores have not been going out on schedule as was originally planned, mainly due to your author getting a lot more work than was originally planned ! Here are the results to date, up to Round 6.

Hopefully I can keep the records a lot more current for the remainder of the competition. Again, remember that the results are also posted on the DURC website : http://www.tcd.ie/clubs/rifle

Mise le meas,

Dublin University Rifle Club ·Trinity College Dublin ·Dublin 2 ·Ireland

Agnes Graham 114 Towerview Crescent Bangor Co. Down BT19 6A2

Mark Dennehy,

DURC Match Secretary

☎: +353-87-2258207

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2000, current scores  9/12/2000

A Chara,

TTT scores have not been going out on schedule as was originally planned, mainly due to your author getting a lot more work than was originally planned ! Here are the results to date, up to Round 6.

Hopefully I can keep the records a lot more current for the remainder of the competition. Again, remember that the results are also posted on the DURC website : http://www.tcd.ie/clubs/rifle

Mise le meas,

Dublin University Rifle Club ·Trinity College Dublin ·Dublin 2 ·Ireland

Ivor Fletcher 127 Belfast Road Antrim Co. Antrim BT41 2AZ

Mark Dennehy,

DURC Match Secretary

☎: +353-87-2258207

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2000, current scores  9/12/2000

A Chara,

TTT scores have not been going out on schedule as was originally planned, mainly due to your author getting a lot more work than was originally planned ! Here are the results to date, up to Round 6.

Hopefully I can keep the records a lot more current for the remainder of the competition. Again, remember that the results are also posted on the DURC website : http://www.tcd.ie/clubs/rifle

Mise le meas,

Dublin University Rifle Club ·Trinity College Dublin ·Dublin 2 ·Ireland

Karen Farrell 41 Deerpark Walk Blessington Co. Wicklow

Mark Dennehy,

DURC Match Secretary

☎: +353-87-2258207

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2000, current scores  9/12/2000

A Chara,

TTT scores have not been going out on schedule as was originally planned, mainly due to your author getting a lot more work than was originally planned ! Here are the results to date, up to Round 6.

Hopefully I can keep the records a lot more current for the remainder of the competition. Again, remember that the results are also posted on the DURC website : http://www.tcd.ie/clubs/rifle

Mise le meas,